The table below shows the salary for two vice presidents of the United States and also the value of the CPI during a year in which each was vice president. The base year of the CPl is an average of 1982 to 1984 dollars. Place your final answer to the question in SNoodle below and show your work on your scratch paper CPI (Base Year 1982-1954) Year Vice President Salay 191 Abn Barkky 2001 $35,000 026 Dick Chency $175,400 What is Alben Barkleys salary in terms of 1982-1984 dollars? What is Dick Cheneys salary in terms of 1982-1984 dollars?
Answer #1

Alben Barkley's salary in terms of 1982-1984 dollars = 35000/0.26 = $134615.39

D. Cheney salary in terms of 1982-1984 dollars = 175400/1.771 = $98814.23